# Benchmarking Energy Calculations Using Formal Proofs

### v1

@@ Abstract
Traditional approaches for validating molecular simulations rely on making software open source and transparent, incorporating unit testing, and generally employing human oversight. We propose an approach that eliminates software errors using formal logic, providing proofs of correctness. We use the Lean theorem prover and programming language to create a rigorous, mathematically verified framework for computing molecular interaction energies. We demonstrate this in LeanLJ, a package of functions, proofs, and code execution software that implements Lennard-Jones energy calculations in periodic boundaries. We introduce a strategy that uses polymorphic functions and typeclasses to bridge formal proofs (about idealized Real numbers) and executable programs (over floating point numbers). Execution of LeanLJ matches the current gold standard NIST benchmarks, while providing even stronger guarantees, given LeanLJ’s grounding in formal mathematics. This approach can be extended to formally verified molecular simulations, in particular, and formally verified scientific computing software, in general.
@@ 1.   Introduction
Molecular simulations serve as a cornerstone for understanding the behaviour of matter at the atomic scale, enabling predictions of properties and phenomena that underpin disciplines ranging from material science to biophysics. Central to these simulations is the accurate modelling of interactions between particles. Among the most widely used models are the Lennard-Jones potential and the Coulomb interaction, which capture the essential physics of van der Waals forces and electrostatic interactions [1] [2] [3] [4] [5]. The Lennard-Jones potential describes non-bonded interactions between particles, particularly in simple fluids and gases [6] [7]. It describes the interaction between a pair of neutral atoms or molecules based on their distance [8] [9], and efficiently captures the balance between attractive and repulsive forces [1] [10].
Practical simulations also incorporate periodic boundary conditions (PBC) to approximate an infinite system by wrapping particles around the edges of the simulation box [11] [12]. This ensures that no particle is artificially confined by the box itself. When computing interactions, the minimum image convention is applied, which means that each particle interacts only with the closest periodic copy of another particle [13]. Using the shortest physically meaningful (periodic) distance between particles avoids redundant calculations and reduces computational overhead. Long-range effects are approximately included via tail corrections.
@@
Software tools like LAMMPS and Gromacs [14][15][16] allow users to simulate the dynamics of large molecular systems. However, the sheer complexity of these software packages and the systems they intend to model presents challenges in making simulations transparent, reproducible, useable by others and extensible (TRUE) [17]. For example, the SAMPL Challenges (Statistical Assessment of the Modelling of Proteins and Ligands) [18] and the Industrial Fluid Properties Simulation Challenges [19][20] task computational researchers to predict the solvation or binding free energies of small molecules or the thermophysical properties of fluids. Each year, researchers submit highly variable answers, reflecting differences in modelling choices by the researchers (e.g. force fields, simulation conditions, free-energy extrapolation strategies, etc.), as well as more hidden, subtle differences amongst software packages (e.g. default settings for managing Lennard-Jones cut-off and settings for Ewald summation). Projects such as the Molecular Simulation Design Framework (MoSDeF) [17][21] and the Molecular Sciences Software Institute (MolSSI) [22] address these issues by providing reproducible workflows for molecular simulation setup, and by teaching and promoting best practices in software development [23]. Simulation software can also be validated by comparing to benchmarks, such as those on the National Institute of Standards and Technology (NIST) Standard Reference Simulation Website (SRSW) [24].
@@
We propose an alternative paradigm for improving reliability of molecular simulations. To illustrate, consider the taxonomy of programming errors in Table 1. The simplest are syntax errors: these are addressed immediately because the code cannot compile, the editor highlights the mistake, and the programmer fixes it. Runtime errors occur during code execution, and may arise when users run the program under conditions not anticipated by the software developers. Nonetheless, runtime errors typically provide a helpful error message pointing toward the source of the issue. The deepest issues are semantic errors in the meaning of the software: Python won’t complain about misinterpreting a scientific principle or incorrectly transcribing math into code – it’s simply not designed for that. Floating-point and round-off errors create numerical inaccuracies, since computers do not operate with infinite mathematical precision. These are addressed by judicious choices of simulation settings and algorithm choices, and by checking conditions like energy conservation after simulation completion [25].
In this work, we propose a strategy for catching syntax, runtime, and semantic errors at the “editor” stage, namely, before the code is compiled. Our approach stems from the formal methods community in computer science, which seeks to prove when software is correct by construction before it is run (also known as static program analysis), unlike traditional testing, which checks for errors by running a program with different inputs. This approach is handy in areas where even small errors can have significant consequences, such as hardware design and critical software systems. A prominent example is the Pentium FDIV bug in Intel processors in the early 1990s, the subject of a multi-million dollar recall stemming from a few misplaced bits in chip software [26]. Now, formal verification approaches prove the correctness of such arithmetic operations in manufactured chips [27]. Our approach most closely resembles that of Selsam, et al., who explored how formal methods can be applied to machine learning systems in Certigrad [28]. By proving the correctness of each step mathematically, this approach exposes errors that might otherwise slip through traditional empirical testing. They highlight the ability of theorem provers like Lean to eliminate entire classes of high-level errors that arise in complex software systems by enforcing correctness through formal reasoning. They demonstrate their approach by building a variational auto-encoder in Lean, proving properties about their implementation of stochastic gradient descent.
@@ 3.   Informal Description of the Molecular Simulation System
Periodic boundary conditions (PBCs) are applied to simulate an infinite system as shown in the equation for particle coordinates in the x, y, and z axes, respectively.
@@
[…]
@@ 4.1.   Introduction to Lean syntax
Here are a few examples to illustrate the syntax of Lean 4. Lean’s basic objects include types, tactics, definitions, and theorems; we don’t introduce any custom types or tactics in this work, so we will focus on definitions and theorems. A definition has the following basic structure [^1] This overview is inspired by the presentation in [35].:
@@ 4.2.   Lennard Jones Potential
Lean allows us to formally prove the equivalence of these two forms, as shown in the theorem lj_eq, allowing us to use either representations confidently. This capability enables not only correctness, but also flexibility in implementing the most efficient forms for simulation. Keep in mind that we don’t address floating-point or round-off errors; this guarantee holds only for idealized functions over Real numbers, which have infinite precision. If a more-efficient version of a function is mathematically equivalent (over Reals) to a base case, but leads to more round-off errors, that won’t be detected in our formulation. A more-efficient version that is not mathematically equivalent (e.g. it invokes an approximation) would be shown to be distinct by this approach.
@@
We can also prove various mathematical properties of our LJ function. Theorem cutoff_behavior states that for any $r>r_{c}$, the value of the Lennard-Jones potential is zero. (The way to read this theorem, is “for Real numbers $\epsilon$, $\sigma$, $r$, and $r_{c}$, assuming $r>r_{c}$, this function evaluates to zero”). This reflects practice of truncating the potential beyond the cut-off distance.
@@ 4.3.   Periodic Boundaries
[…]
@@ 4.4.   Minimum Image Distance
We can prove a neat property of how the periodic boundaries interact with the minimum image distance – that the minimum image distance between arbitrary points in space is equivalent to the minimum image distance between those points, after being wrapped into the simulation box. This is stated in theorem squaredminImageDistance_theorem, which requires an inline invoking a $\lambda$ function to iterate over the box dimensions (for brevity, the proof steps are omitted here, but available on GitHub). This only holds for nonzero box lengths.
@@ 5.1.   Functional Programming
Here, a recursive function energy is deined locally, as well an accumulation variable acc. In the central function call, energy (i+1) j (acc + lj_Float r r_c $\epsilon$ $\sigma$), energy adds one LJ energy contribution to the value of acc, using particle indices $i+1$ and $j$ to obtain the distance $r$. The remaining conditions handle increments on the edge cases.
However, we found neither of these functions to be particularly amenable to proofs. We are particularly looking for a proof that the total number of pairs in the system is $N*(N-1)/2$. To do so, we defined a helper function that generates the pairs. Our function pairs returns a list of pairs of atom indexes (e.g. [(0,1), (0,2), (1,2)] when $n=3$).
@@
We prove the length of this list is $N*(N-1)/2$ in pairs_length_eq in a long proof that uses mathematical induction (full proof available on LeanLJ GitHub). In the absence of test data, such a proof could provide confidence that the implementation is correct. We anticipate that such proofs would be useful for validating energy calculations over molecules with complex exclusion rules (like excluding adjacent up to 1-4 interactions in molecules), and higher-order interactions, like 3-body potentials.
@@
Lean automatically checks functions for termination, which is quite important for recursive functions, lest they get trapped in an infinite loop. In developing this code, we found execution and comparing to the NIST tests was valuable for developing the logic of the loops and the recursion, as it enabled quick feedback. Our code also incorporates tail recursion to facilitate efficient execution [33].
@@ 5.2.   Polymorphism
Figure 6: Explanation of the polymorphic pbc function. The function is defined over a generic type $\alpha$, and the required operations—subtraction, multiplication, division, and rounding—are expressed through typeclasses: HSub, HMul, HDiv, and HasRound. Each typeclass specifies that the type $\alpha$ must support a given operation. For example, HSub $\alpha$ $\alpha$ $\alpha$ means $\alpha$ must support subtraction with two $\alpha$ inputs returning an $\alpha$ result. Concrete instances, such as Float and $\mathbb{R}$, implement these typeclasses to enable polymorphic behavior. This allows pbc to work with different numeric types, as long as they satisfy the required operations.
@@
To illustrate, consider the function pbc (Fig. 6), which wraps a particle’s position into the simulation box using periodic boundary conditions. Section 4 showed pbc_Real, which operates on position and box length with type $\mathbb{R}$; pbc operates on position and box length with generic type $\alpha$. We tell Lean more about what $\alpha$ can be, by using typeclasses and instances. Specially, pbc is defined for any type $\alpha$ that “knows how to” subtract, multiply, divide, and round. These capabilities are provided through the typeclasses HSub, HMul, HDiv, and HasRound. For example, the typeclass HSub [$\alpha$ $\alpha$ $\alpha$] requires that there exists a definition of subtraction between two members of $\alpha$ that would output a third member of $\alpha$. HSub, HMul, and HDiv are all defined in Mathlib; for rounding, we defined a custom typeclass, since Mathlib didn’t already define that connection. This approach allows our definition of pbc to be used in two very different ways: with Real numbers for formal proofs, and with floating-point numbers for actual simulations.
@@
Prefacing each function with a long list of instances is rather cumbersome. We can group all of these instances together into a single typeclass, RealLike, named thus because it’s “like” Real numbers. RealLike is ultimately compatible with Floats, but it doesn’t have all the properties of the Real numbers (after all, it cannot use functions we didn’t attach to the typeclass). The RealLike typeclass captures the essential features of types that behave like real numbers, making it easier to write numeric code that works across different types such as Float or Int. It includes basic operations like addition, subtraction, multiplication, division, negation, and exponentiation with natural numbers that appear in our equations (1, 3, 9, etc.). It also provides comparisons ($\leq$, $<$) that return Bool, along with constants like zero and one, and ways to convert from natural and integer literals. In addition, RealLike extends two smaller typeclasses—HasSqrt for square roots (used in distance calculations) and HasRound for rounding (used in periodic boundaries) so that any type marked as RealLike is also expected to support those operations. This setup allows us to write clean, reusable, and type-safe numeric functions without tying them to one specific number type.
@@ 6.   Results
30 $-1.67903\mathrm{E}{+01}$ $-1.6790\mathrm{E}{+01}$ $-5.45166\mathrm{E}{-01}$ $-5.4517\mathrm{E}{-01}$
200 $-6.90004\mathrm{E}{+02}$ $-6.9000\mathrm{E}{+02}$ $-2.42296\mathrm{E}{+01}$ $-2.42296\mathrm{E}{+01}$
400 $-1.14666\mathrm{E}{+03}$ $-1.1467\mathrm{E}{+03}$ $-4.96222\mathrm{E}{+01}$ $-4.9622\mathrm{E}{+01}$
800 $-4.35154\mathrm{E}{+03}$ $-4.3515\mathrm{E}{+03}$ $-1.98488\mathrm{E}{+02}$ $-1.9849\mathrm{E}{+02}$
7.   Conclusions and Outlook

### v2

@@ Abstract
Traditional approaches for validating molecular simulations rely on making software open source and transparent, incorporating unit testing, and generally employing human oversight. We propose an approach that eliminates software errors using formal logic, providing proofs of correctness. We use the Lean theorem prover and programming language to create a rigorous, mathematically verified framework for computing molecular interaction energies. We demonstrate this in LeanLJ, a package of functions, proofs, and code execution software that implements Lennard-Jones energy calculations in periodic boundaries. We introduce a strategy that uses polymorphic functions and type classes to bridge formal proofs (about idealized Real numbers) and executable programs (over floating point numbers). Execution of LeanLJ matches the current gold standard NIST benchmarks, while providing even stronger guarantees, given LeanLJ’s grounding in formal mathematics. This approach can be extended to formally verified molecular simulations in particular and formally verified scientific computing software, in general.
@@ 1.   Introduction
Molecular simulations constitute an essential computational tool for understanding atomic-scale phenomena, playing a critical role in predicting the physicochemical properties underlying fields such as materials science, chemistry, and biophysics [1]. For instance, accurate simulations enable the study of molecular interactions that dictate gas adsorption behaviours in porous materials [2] [3], influence solvation dynamics in chemical solutions [4], and affect catalytic reactions on material surfaces [5]. Accurate modelling of particle interactions lies at the core of molecular simulations [6] . Among the most fundamental and extensively used models are the Lennard-Jones potential, describing van der Waals interactions, and Coulomb potentials, capturing electrostatic forces. The Lennard-Jones potential, specifically, finds significant application in representing non-bonded interactions in chemically relevant systems such as simple fluids [7], noble gas clusters [8], hydrocarbon fluids [9], and molecular adsorption phenomena in zeolites or metal-organic frameworks (MOFs) [10] [11]. It also describes the interaction between a pair of neutral atoms or molecules based on their distance [12] [13] [14], and efficiently captures the balance between attractive and repulsive forces [15] [16]. Coulomb interactions are particularly critical for describing electrolyte solutions [17], protein-ligand binding affinities [18], and charged colloidal systems [19] [18].
To accurately approximate infinite molecular systems from computationally manageable finite-sized simulations, periodic boundary conditions (PBC) are conventionally employed [20] [21]. In chemical simulations, PBC enables the study of bulk-phase properties without boundary effects, such as predicting phase behaviours in liquid water [22], ionic solutions [23], or polymer melts [24]. Furthermore, the minimum image convention ensures that computational resources are efficiently utilized by considering only the nearest periodic images in calculations [25], which is particularly important in simulations of dense chemical environments like ionic liquids or liquid crystal phases.
@@
Software tools like LAMMPS and Gromacs [26][27][28] allow users to simulate the dynamics of large molecular systems. However, the sheer complexity of these software packages and the systems they intend to model presents challenges in making simulations transparent, reproducible, useable by others and extensible (TRUE) [29]. For example, the SAMPL Challenges (Statistical Assessment of the Modelling of Proteins and Ligands) [30] and the Industrial Fluid Properties Simulation Challenges [31][32] task computational researchers to predict the solvation or binding free energies of small molecules or the thermophysical properties of fluids. Each year, researchers submit highly variable answers, reflecting differences in modelling choices by the researchers (e.g. force fields, simulation conditions, free-energy extrapolation strategies, etc.), as well as more hidden, subtle differences amongst software packages (e.g. default settings for managing Lennard-Jones cut-off and settings for Ewald summation). Projects such as the Molecular Simulation Design Framework (MoSDeF) [33] [29][34] and the Molecular Sciences Software Institute (MolSSI) [35] address these issues by providing reproducible workflows for molecular simulation setup, and by teaching and promoting best practices in software development [36]. Simulation software can also be validated by comparing to benchmarks, such as those on the National Institute of Standards and Technology (NIST) Standard Reference Simulation Website (SRSW) [37].
@@
We propose an alternative paradigm for improving reliability of molecular simulations. To illustrate, consider the taxonomy of programming errors in Table 1. The simplest are syntax errors: these are addressed immediately because the code cannot compile, the editor highlights the mistake, and the programmer fixes it. Runtime errors occur during code execution, and may arise when users run the program under conditions not anticipated by the software developers. Nonetheless, runtime errors typically provide a helpful error message pointing toward the source of the issue. The deepest issues are semantic errors in the meaning of the software: Python would not complain about misinterpreting a scientific principle or incorrectly transcribing math into code – it is simply not designed for that. Floating-point and round-off errors create numerical inaccuracies, since computers do not operate with infinite mathematical precision. These are addressed by judicious choices of simulation settings and algorithm choices, and by checking conditions like energy conservation after simulation completion [38].
In this work, we propose a strategy for catching syntax, runtime, and semantic errors at the “editor” stage, namely, before the code is compiled. Our approach stems from the formal methods community in computer science, which seeks to prove when software is correct by construction before it is run (also known as static program analysis), unlike traditional testing, which checks for errors by running a program with different inputs. This approach is handy in areas where even small errors can have significant consequences, such as hardware design and critical software systems. A prominent example is the Pentium FDIV bug in Intel processors in the early 1990s, the subject of a multi-million dollar recall stemming from a few misplaced bits in chip software [39]. Now, formal verification approaches prove the correctness of such arithmetic operations in manufactured chips [40]. Our approach most closely resembles that of Selsam, et al., who explored how formal methods can be applied to machine learning systems in Certigrad [41] (Figure 1). By proving the correctness of each step mathematically, this approach exposes errors that might otherwise slip through traditional empirical testing. They highlight the ability of theorem provers like Lean to eliminate entire classes of high-level errors that arise in complex software systems by enforcing correctness through formal reasoning. They demonstrate their approach by building a variational auto-encoder in Lean, proving properties about their implementation of stochastic gradient descent.
@@ 3.   Informal Description of the Molecular Simulation System
Periodic boundary conditions (PBCs) are applied to simulate an infinite system as shown in the equation for particle coordinates in the x, y, and z axes, respectively (Figure 2).
@@
$U_{LRC}$ $=\frac{1}{2}4\pi\rho\int^{\infty}_{r_{c}}r^{2}V_{LJ}(r)dr$ (9)
$=\frac{1}{2}4\pi\rho\int^{\infty}_{r_{c}}r^{2}4\epsilon\left(\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right)dr$ (10)
@@ 4.1. Introduction to Lean syntax
Here are a few examples to illustrate the syntax of Lean 4. Lean’s basic objects include types, tactics, definitions, and theorems; we do not introduce any custom types or tactics in this work, so we will focus on definitions and theorems. A definition has the following basic structure [^1] This overview is inspired by the presentation in [48].:
@@ 4.2. Lennard Jones Potential
Lean allows us to formally prove the equivalence of these two forms, as shown in the theorem lj_eq, allowing us to use either representations confidently. This capability enables not only correctness, but also flexibility in implementing the most efficient forms for simulation. Keep in mind that we do not address floating-point or round-off errors; this guarantee holds only for idealized functions over Real numbers, which have infinite precision. If a more-efficient version of a function is mathematically equivalent (over Reals) to a base case, but leads to more round-off errors, that would not be detected in our formulation. A more efficient version that is not mathematically equivalent (e.g. it invokes an approximation) would be shown to be distinct by this approach.
@@
We can also prove various mathematical properties of our LJ function. Theorem cutoff_behaviour states that for any $r>r_{c}$, the value of the Lennard-Jones potential is zero. (The way to read this theorem, is “for Real numbers $\epsilon$, $\sigma$, $r$, and $r_{c}$, assuming $r>r_{c}$, this function evaluates to zero”). This reflects the practice of truncating the potential beyond the cut-off distance.
@@ 4.3. Periodic Boundaries
We formally proved that the wrapped displacement produced by the periodic boundary condition function lies within the interval [L/2,L/2][-L/2,L/2] for any Real coordinate $p$ and positive box length $L$. This ensures that particles always interact with the nearest periodic image, which is a key assumption in molecular dynamics simulations. The proof was constructed in Lean by expressing the wrapped position as $L\cdot\delta$, where $\delta=\frac{p}{L}-\text{round}\left(\frac{p}{L}\right)$, and rigorously showing that $|\delta|\leq\frac{1}{2}$, hence $|\text{pbc\_Real}(p,L)|\leq\frac{L}{2}$.
@@ 4.4. Minimum Image Distance
We can prove a neat property of how the periodic boundaries interact with the minimum image distance – that the minimum image distance between arbitrary points in space is equivalent to the minimum image distance between those points, after being wrapped into the simulation box. This is stated in theorem squaredminImageDistance_theorem, which requires an inline invoking a $\lambda$ function to iterate over the box dimensions (for brevity, the proof steps are omitted here, but available on GitHub). This only holds for non-zero box lengths.
@@ 5.1. Functional Programming
Here, a recursive function energy is defined locally, as well an accumulation variable acc. In the central function call, energy (i+1) j (acc + lj_Float r r_c $\epsilon$ $\sigma$), energy adds one LJ energy contribution to the value of acc, using particle indices $i+1$ and $j$ to obtain the distance $r$. The remaining conditions handle increments on the edge cases.
However, we found neither of these functions to be particularly amenable to proofs. We are particularly looking for a proof that the total number of pairs in the system is N(N1)/2N*(N-1)/2. To do so, we defined a helper function that generates the pairs. Our function pairs returns a list of pairs of atom indexes (e.g. [(0,1), (0,2), (1,2)] when $n=3$).
@@
We prove the length of this list is N(N1)/2N*(N-1)/2 in pairs_length_eq in a long proof that uses mathematical induction (full proof available on LeanLJ GitHub). In the absence of test data, such a proof could provide confidence that the implementation is correct. We anticipate that such proofs would be useful for validating energy calculations over molecules with complex exclusion rules (like excluding adjacent up to 1-4 interactions in molecules), and higher-order interactions, like 3-body potentials.
@@
Lean automatically checks functions for termination, which is quite important for recursive functions, lest they get trapped in an infinite loop. These are verified formally using tactics that operate behind-the-scenes; only in more complicated cases, in which Lean cannot infer termination automatically, will the user be prompted to write a termination proof. Likewise, Lean also ensures that array indexing is safe, preventing all runtime errors associated with accessing out-of-bounds indices. In developing this code, we found execution and comparing to the NIST tests was valuable for developing the logic of the loops and the recursion, as it enabled quick feedback. Our code also incorporates tail recursion to facilitate efficient execution [46].
@@ 5.2. Polymorphism
Figure 6: Explanation of the polymorphic pbc function. The function is defined over a generic type $\alpha$, and the required operations—subtraction, multiplication, division, and rounding—are expressed through type classes: HSub, HMul, HDiv, and HasRound. Each type class specifies that the type $\alpha$ must support a given operation. For example, HSub $\alpha$ $\alpha$ $\alpha$ means $\alpha$ must support subtraction with two $\alpha$ inputs returning an $\alpha$ result. Concrete instances, such as Float and $\mathbb{R}$, implement these type classes to enable polymorphic behaviour. This allows pbc to work with different numeric types, as long as they satisfy the required operations.
@@
To illustrate, consider the function pbc (Fig. 6), which wraps a particle’s position into the simulation box using periodic boundary conditions. Section 4 showed pbc_Real, which operates on position and box length with type $\mathbb{R}$; pbc operates on position and box length with generic type $\alpha$. We tell Lean more about what $\alpha$ can be, using type classes and instances. Specially, pbc is defined for any type $\alpha$ that “knows how to” subtract, multiply, divide, and round. These capabilities are provided through the type classes HSub, HMul, HDiv, and HasRound. For example, the type class HSub $\alpha$ $\alpha$ $\alpha$] requires that there exists a definition of subtraction between two members of $\alpha$ that would output a third member of $\alpha$. HSub, HMul, and HDiv are all defined in Mathlib; for rounding, we defined a custom type class, since Mathlib did not already define that connection. This approach allows our definition of pbc to be used in two very different ways: with Real numbers for formal proofs, and with floating-point numbers for actual simulations.
@@
Prefacing each function with a long list of instances is rather cumbersome. We can group all of these instances together into a single type class, RealLike, named thus because it is “like” Real numbers. RealLike is ultimately compatible with Floats, but it doesn’t have all the properties of the Real numbers (after all, it cannot use functions we did not attach to the type class). The RealLike type class captures the essential features of types that behave like real numbers, making it easier to write numeric code that works across different types such as Float or Int. It includes basic operations like addition, subtraction, multiplication, division, negation, and exponentiation with natural numbers that appear in our equations (1, 3, 9, etc.). It also provides comparisons ($\leq$, <<) that return Bool, along with constants like zero and one, and ways to convert from natural and integer literals. In addition, RealLike extends two smaller type classes—HasSqrt for square roots (used in distance calculations) and HasRound for rounding (used in periodic boundaries) so that any type marked as RealLike is also expected to support those operations. This setup allows us to write clean, reusable, and type-safe numeric functions without tying them to one specific number type.
We note that this type class approach admits a small possibility of error that Lean will not catch. The RealLike type class manually links certain Real-typed and Float-typed functions, and mistakes in RealLike won’t be flagged by Lean. For example, a subtle error in which one rounding function rounded 0.5 down while another rounded 0.5 up would not be detected by Lean. In fact, egregious errors linking a square root and a cube root are also technically possible, so human oversight remains necessary with this approach. Essentially, the RealLike type class defines which Real and Float functions are semantically equivalent, and the human is responsible for ensuring correct semantics. The Lean community is developing more systematic solutions, such as the ComputableReal project led by Alex Meiburg, which provides a more principled, computable representation of real numbers that integrates cleanly with existing numeric type classes [53].
@@ 6.   Results
30 $1.67903\mathrm{E}{+01}$ $1.6790\mathrm{E}{+01}$ $5.45166\mathrm{E}{-01}$ $5.4517\mathrm{E}{-01}$
200 $6.90004\mathrm{E}{+02}$ $6.9000\mathrm{E}{+02}$ $2.42296\mathrm{E}{+01}$ $2.42296\mathrm{E}{+01}$
400 $14666\mathrm{E}{+03} 1.1$$1467\mathrm{E}{+03}$ $4.96222\mathrm{E}{+01}$ $4.9622\mathrm{E}{+01}$
800 $4.35154\mathrm{E}{+03}$ $4.3515\mathrm{E}{+03}$ $1.98488\mathrm{E}{+02}$ $1.9849\mathrm{E}{+02}$
7.   Discussion and Outlook